Find the inverse function of 𝑓(𝑥) = −𝑥2 + 8𝑥 − 14 if 𝑥 ≤ 4. Find the domain of this inverse function.


1
Expert's answer
2021-11-01T17:45:12-0400
"f(x)-x^2+8x-14=-(x^2-8x+16)+2"

"=-(x-4)^2+2"

The function "f(x)=-x^2+8x-14, x\leq4" is one-to-one. Then it has inverse.

Replace "f(x)" with "y"


"y=-(x-4)^2+2, x\leq4, y\leq2"

 Replace every "x" with a "y" and replace every "y"  with an "x"

"x=-(y-4)^2+2, x\leq2, y\leq4"

Solve for "y"


"(y-4)^2=2-x, x\leq2, y\leq4"

"y-4=-\sqrt{2-x}"

"y=4-\sqrt{2-x}"

Replace "y" with "f^{-1}(x)"


"f^{-1}(x)=4-\sqrt{2-x}"

Domain "(-\infin, 2]"